1 /xy + y is an integrating factor for the differential equation (x^2y + y^2)dx +( y^3-x^3)dy=0. True or false an


1
Expert's answer
2021-03-24T14:35:53-0400

Solution

Let P(x,y) = x2y + y2, Q(x,y) =  y3-x3, M(x,y) =  1 /xy + y.

If M(x,y) is an integrating factor then next condition is to be satisfied:

∂(M*P)/∂y = ∂(M*Q)/∂x

M*P = (x2y + y2)*(1 /xy + y) = x+y/x+x2y2+y3  =>  ∂(M*P)/∂y = 1/x+2x2y+3y2 

M*Q = (y3-x3)* (1 /xy + y) = y2/x-x2/y+y4-x3y  =>  ∂(M*Q)/∂x = -y2/x2-2x/y-3x2y

Therefore ∂(M*P)/∂y ≠ ∂(M*Q)/∂x.

So it’s False that M(x,y) =  1 /xy + y is an integrating factor for the differential equation (x2y + y2)dx +(y3-x3)dy=0.

Answer

False